Need assistance ASAP!!! Got the adjusted trial balance already
Use this page for tow Crestwood Adjusted Trial Balance December 31, 20X1 Debit $17.170 Credis Cash A/R ADA Pre-paid insurance 16,900 2,028 Building 170,000 31,2 Accumulated Depn-Building Accumulated Salary payable Common stock Retained carnings Dues revenue 08, 140 9S,210 21,350 Maintenance expense Bad debt expense Uncarned revenue Rent receivable Depreciation expense-building Depreciation expense-equipment Insurance expense 0 4000 37 30
Crestwood Income Statement For Year Ended December 31, 20X1 Revenues: Expenses: Net Income
Crestwood Statement of Retained Earnings For Year Ended December 31, 20x1 Beginning R/E Ending R/E s.
Crestwood Balance Sheet As of December 31, 20X1 Assets: Total Current Assets $ Total Assets Liabilities and Owners Equity Total Liabilities S Total Equity Total Liabilities and Owners Equity $
Post Closing Trial Balance As of December 31, 20x1 Credin Debit Use this page to record closing entries. Close out revenues Close out expenses. Close out Income Summary account.
Post Closing Trial Balance As of December 31, 20xI Ascount Cash A/R ADA Pre-paid insurance Land Building Accumulated Depn-Building Equip Accumulated Depn-Equipment Salary payable Common stock Retained earnings Debit $17,170 Crodit 16,900 Unearned dues revenue Rent receivable 10
Answer #1
A B C D E F G H
2
3
4
5 Adjusting Entries
6 Account Code Account Adjusted Trial Balance
7 Debit Credit
8 1 Cash $17,170
9 2 Accounts Receivable $16,900
10 3 ADA (Allowance for doubtful accounts) $2,028
11 4 Prepaid Insurance $5,240
12 7 Land $358,560
13 8 Building $170,000
14 9 Accumulated Dep. - Building $39,260
15 10 Equip. $122,340
16 11 Accumulated Dep. - Equip $89,114
17 12 Salary Payable $3,530
18 13 Common Stock $408,140
19 14 Retained Earnings $119,040
20 15 Dues Revenue $191,160
21 16 Greens fees $6,400
22 17 Rent Revenue $19,200
23 18 Utilities Expense $56,430
24 19 Salary Expense $85,220
25 20 Maintenance Expense $27,350
26 21 Bad Debt Expense $998
27 22 Unearned Revenue $9,000
28 23 Rent receivable $1,600
29 24 Depreciation Exp - Building $4,000
30 25 Depreciation Exp - Equipment $17,234
31 26 Insurance Expense $3,830
32 Total $886,872 $886,872
33
34
35 Income Statement
36 Dues Revenue $191,160.00 =F20
37 Greens fees $6,400.00 =F21
38 Rent Revenue $19,200.00 =F22
39 Total Revenue $216,760.00 =SUM(E36:E38)
40 Expenses:
41 Utilities Expense $56,430.00 =E23
42 Salary Expense $85,220.00 =E24
43 Maintenance Expense $27,350.00 =E25
44 Bad Debt Expense $998.00 =E26
45 Depreciation Exp - Building $4,000.00 =E29
46 Depreciation Exp - Equipment $17,234.00 =E30
47 Insurance Expense $3,830.00 =E31
48 Total Expense $195,062.00 =SUM(E41:E47)
49 Net Income $21,698.00 =E39-E48
50
51 Statement of Retained Earnings
52 Beginning Retained Earnings $119,040.00 =F19
53 Add: Net Income $21,698.00 =E49
54 Less: Dividends $0.00 =E19
55 Ending Retained Earnings $140,738.00 =E52+E53-E54
56
57 Balance Sheet
58 Cash $17,170.00 =E8
59 Accounts Receivable $16,900.00 =E9
60 ADA (Allowance for doubtful accounts) ($2,028.00) =-F10
61 Prepaid Insurance $5,240.00 =E11
62 Rent Receivable $1,600.00 =E28
63 Current Assets $38,882.00 =SUM(E58:E62)
64 Land $358,560.00 =E12
65 Building $170,000.00 =E13
66 Accumulated Dep. - Building ($39,260.00) =-F14
67 Equip. $122,340.00 =E15
68 Accumulated Dep. - Equip ($89,114.00) =-F16
69 Total Assets $561,408.00 =E63+SUM(E64:E68)
70
71 Liabilities and Shareholder's Equity
72 Salaries Payable $3,530.00 =F17
73 Unearned commission revenue $9,000.00 =F27
74 Total Liabilities $12,530.00 =SUM(E72:E73)
75 Common Stock $408,140.00 =F18
76 Retained Earnings $140,738.00 =E55
77 Total Liabilities and Shareholder's Equity $561,408.00 =E74+E75+E76
78